OpenStudy (anonymous):

In the legend of the King and inventor of the chessboard, how many grains of wheat were on the 9th square of the chessboard? A. 2 B. 9 C. 256 D. 512

OpenStudy (anonymous):

Hi Cydnim, I am Mikael and think of each next-square as multiplying the previous number of grains by 2

OpenStudy (anonymous):

Sooo Square 1 is \[2^{0}\]

OpenStudy (anonymous):

So it would be 2 ?

OpenStudy (anonymous):

Square 2 is \[2^{1}\]

OpenStudy (compassionate):

Wrong section, And 256

OpenStudy (anonymous):

Square 3 is\[2^{?}\]

OpenStudy (anonymous):

Anyway square N is \[2^{N-1}\]

OpenStudy (anonymous):

Sooo ... if we consider Square N= 9 then it has 2^(9-1) = 2^8 = 256

OpenStudy (anonymous):

Thank you so much !! (:

OpenStudy (anonymous):

Medals Cydnim pls, Fan is optional